if a box has 2.0x10^2J of work done to it, is moved with 55N. what is the distance it moved?


1
Expert's answer
2021-03-21T11:23:42-0400

By the definition of the work done, we have:

"W=Fd,""d=\dfrac{W}{F}=\dfrac{2.0\cdot10^2\ J}{55\ N}=3.64\ m."